# On modules whose dual is of finite Gorenstein dimension: The $k$-torsion, homological criteria, and applications in modules of differentials of order $n$

### v4

@@ 4. Characterization of totally reflexives
Let $R$ be a local ring of depth $t$ and let $M$ be a non-zero $R$ module. Suppose that $M$ is locally totally reflexive on the punctured spectrum of $R$. Suppose that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following are equivalent:
@@
The equivalences $(1)\Leftrightarrow(2)\Leftrightarrow(3)$ follow from Theorem 4.9, while $(3)\Leftrightarrow(1)$ follow from Theorem 4.1. ∎
@@ 5. Freeness Criteria
    To discuss the Auslander-Reiten conjecture for $R$-modules such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$.
@@
Let $R$ be a Cohen-Macaulay local ring and let $M$ be an $R$-module such that $\operatorname{pd}_{R}(M^{\ast})<\infty$. If $M$ is maximal Cohen-Macaulay, then $M$ is free.
Proof.
This follows immediately from Theorem 5.1 by taking $n=\operatorname{dim}(R)$. ∎
@@
Let $R$ be a Cohen-Macaulay local ring and let $M$ be an $R$-module such that $M^{\ast}\not=0$ and $\operatorname{pd}_{R}(M^{\ast})<\infty$. If $M$ is Cohen-Macaulay, then $M$ is free.
Proof.
Since $M^{\ast}\not=0$, note that $\operatorname{grade}_{R}(M)=0$. Moreover, by [7, Corollary 2.1.4], $\operatorname{grade}_{R}(M)=\operatorname{dim}_{R}(R)-\operatorname{dim}_{R}(M)$ and hence $\operatorname{dim}_{R}(M)=\operatorname{dim}_{R}(R).$ Thus from Cohen-Macaulayness of $M$, we see that $M$ is maximal Cohen-Macaulay. Hence, it follows from Corollary 5.2 that $M$ is free. ∎
@@
Let $R$ be a local Cohen-Macaulay. The maximal ideal $\mathfrak{m}$ is Cohen-Macaulay if and only $\operatorname{dim}(R)\leq 1$.
@@
Since $\operatorname{depth}_{R}(R)\geq 2,$ then $k^{\ast}=0=\operatorname{Ext}_{R}^{1}(k,R)$ and $\mathfrak{m}^{\ast}\cong R$. Therefore, by Corollary 5.3, $\mathfrak{m}$ is free, and hence $\mathfrak{m}\cong R$. From the exact sequence (5.1), we obtain an exact sequence
@@
As $\operatorname{depth}_{R}(R)\geq 2,$ then $\operatorname{Hom}_{R}(k,\mathfrak{m})\cong\operatorname{Hom}_{R}(k,R)=0=\operatorname{Ext}_{R}^{1}(k,R)\cong\operatorname{Ext}_{R}^{1}(k,\mathfrak{m})$. Consequently, $\operatorname{Hom}_{R}(k,k)=0,$ which does not occurs.
Conversely, assume that $\operatorname{dim}(R)\leq 1$. If $\dim(R)=0$ it is clear that all $R$-modules are Cohen-Macaulay. Suppose $\operatorname{dim}R(R)=1$. Since $R$ is Cohen-Macaulay, then there exists an $R$-regular element $x\in\mathfrak{m}$. Note that $x$ is also an $\mathfrak{m}$-regular element. Therefore $1\leq\operatorname{depth}_{R}(\mathfrak{m})\leq\operatorname{dim}_{R}(R)=1,$ which shows that $\mathfrak{m}$ is maximal Cohen-Macaulay. ∎
@@
Let $R$ be a local ring of depth $t\geq 1$. Let $M$ be a locally free $R$-module on the spectrum punctured such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{t-1}(M,M)=0$ for all $1\leq i\leq t$, then $M$ is free.
@@
If $t=1$, then $\operatorname{Hom}_{R}(M,M)\cong\operatorname{Ext}_{R}^{0}(M,M)=0$, whence $M$ is zero and hence free. Suppose that $t\geq 2$. By Theorem 4.1, $M$ is totally reflexive and hence $\operatorname{Tr}(M)$ as well. Thus, $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $i>0$. Therefore, the result follows from [30, Proposition 2.8]. ∎
@@
Let $R$ be a local ring of depth $t\geq 1$. Let $M$ be a locally free $R$-module on the spectrum punctured such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $0pt(M)\geq t$ and $\operatorname{Ext}_{R}^{t-1}(M,M)=0$, then $M$ is free.
@@
Let $R$ be a ring satisfying $(S_{1})$ and $X$ a subset of $\operatorname{Spec}(R)$ containing $X^{1}(R)$. Let
$s:=\inf\left\{\operatorname{depth}R_{\mathfrak{p}}\mid\mathfrak{p}\in\operatorname{Spec}(R)\backslash X\right\}$ and $t:=\sup\left\{\operatorname{depth}R_{\mathfrak{p}}\mid\mathfrak{p}\in\operatorname{Spec}(R)\backslash X\right\}$.
Suppose that $M$ is locally free on $X$ and that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $s-1\leq j\leq t-1$, then $M$ is projective.
@@
We may suppose that $X\not=\operatorname{Spec}(R)$. We show that $M_{\mathfrak{p}}$ is free for all $\mathfrak{p}\in\operatorname{Spec}(R)$. By assumption this is true if $\mathfrak{p}\in X$. So, suppose that $\mathfrak{p}\in\operatorname{Spec}(R)\backslash X$. Therefore $\operatorname{ht}(\mathfrak{p})\geq 2$ and $\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\geq 1$ since $R$ satisfies $(S_{1})$. By definition of $s$ and $t$, note that $s\leq\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\leq t$, so $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(M_{\mathfrak{p}},R_{\mathfrak{p}})=0$ for all $1\leq i\leq\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$ and $\operatorname{Ext}_{R_{\mathfrak{p}}}^{\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})-1}(M_{\mathfrak{p}},M_{\mathfrak{p}})=0$. Hence, by Proposition 5.5, $M_{\mathfrak{p}}$ is free. ∎
@@
([30, Remark 2.11]) Let $R$ be a Cohen-Macaulay local ring of dimension $d$ and $n$ be an integer such that $11\leq n\leq d-$ 1. Then
@@
$\sup\{\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\mid\mathfrak{p}\in\operatorname{Spec}(R)\backslash\mathrm{X}^{n}(R)\}=d.$
@@
(1) We claim that $M$ is locally free on $X^{n}(R)$. In fact, given $\mathfrak{p}\in X^{n}(R)$, we have by hypothesis that $\operatorname{pd}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})<\infty$ and $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(M_{\mathfrak{p}},R_{\mathfrak{p}})=0$ for all $1\leq i\leq d$. Hence, by [38, p. 154, Lemma 1(iii)], $M_{\mathfrak{p}}$ is free as $R_{\mathfrak{p}}$-module. Thus item (1) follows immediately from Theorem 5.7 and Remark 5.8.
(2) We separate the proof in two cases. Let $p=\operatorname{G-dim}_{R}(M)$. We prove the result by considering two cases.
@@
(2.b) Suppose $p>0$. We have $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i\geq p+1$. By [30, Lemma 2.6], we get
@@
Let $R$ be a Cohen-Macaulay local ring of dimension $d\geq 2$, and let $M$ be a maximal Cohen-Macaulay $R$-module such that $\operatorname{G-dim}_{R}(M)<\infty$. Let $11\leq n\leq d-$ 1 be an integer such that $M$ is locally of finite projective dimension on $X^{n}(R)$. If $\operatorname{Ext}_{R}^{i}(M,M)=0$ for all $n\leq i\leq d-1$. Then $M$ is free.
Proof.
As $\operatorname{G-dim}_{R}(M)<\infty$ and $M$ is maximal Cohen-Macaulay, it follows from the Auslander-Bridger Formula that $\operatorname{G-dim}_{R}(M)=0$. Then, by Theorem 5.9(2), $\operatorname{pd}_{R}(M)<\infty$, and hence $\operatorname{pd}_{R}(M)=\operatorname{G-dim}_{R}(M)=0$. This yields that $M$ is free. ∎
5.11 Corollary.
Let $R$ be a Cohen-Macaulay normal local ring of dimension $d\geq 2$, and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Suppose that $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq d$ and $11\leq j\leq d-$ 1. Then $M$ is free.
Proof.
As $R$ is normal, then $R_{\mathfrak{p}}$ is a regular local ring for all $\mathfrak{p}\in X^{1}(R)$. Hence, $M$ is locally of finite projective dimension on $X^{1}(R)$. Thus, the result follows from Theorem 5.9. ∎
@@
In [12, Corollary 6.9(2)], Dey and Ghosh demonstrated that the Auslander-Reiten conjecture holds for (finitely generated) $R$-modules $M$ satisfying $\operatorname{G-dim}_{R}(M)<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Motivated by this, we investigate whether the conjecture remains valid when we consider $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ instead of $\operatorname{G-dim}_{R}(M)<\infty$ in the referenced case.
@@
    $\operatorname{Hom}_{R}(M,M)$ is free and $\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $11\leq j\leq t-$ 1.
@@
    $\operatorname{Hom}_{R}(M,M)$ has finite projective dimension and $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $11\leq j\leq t-$ 1.
@@
$(2)\Rightarrow(3)$. We only need to prove that $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t$. For this, we will show that $M$ is totally reflexive. In view of [12, Corollary 7.6(2)], $M$ satisfies $(\widetilde{S}_{t})$. Therefore, $M$ is totally reflexive by Theorem 4.1.
$(3)\Rightarrow(1)$. By Theorem 4.1, $M$ is totally reflexive. Therefore by [12, Corollary 6.8], $M$ is free. ∎
@@
Let $R$ be local ring. The Auslander-Reiten conjecture holds true for all (finitely generated) $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$.
@@
Let $R$ be a local ring, and let $C$ be a semidualizing $R$-module. If $\operatorname{G-dim}_{R}(C^{\ast})<\infty$, then $C\cong R$.
@@ 6. Gorenstein criteria and related questions
Throughout this section, let $R$ be a local ring. In this section, we aim to provide a number of criteria for a local ring $R$ to be Gorenstein in terms of the dual of certain modules having finite Gorenstein dimension. In this sense, we discuss some related questions to this subject.
The first criterion that we present in this section is an application of Theorem 4.2. Before stating it, first recall that if $R$ is a ring with total quotient ring $Q$, then an $R$-module $M$ is said to have a (generic) rank, denoted by $\operatorname{rank}(M)$, if $M\otimes_{R}Q$ is a free $Q$-module of rank $\operatorname{rank}(M)$. Let $e(R)$ denote the Hilbert-Samuel multiplicity of $R$ and $\mu(M)$ denote the minimum number of generators of $M$. Moreover, if $M$ is Cohen-Macaulay and $e(M)=\mu(M)$, then $M$ it is said to be an Ulrich module ([21, Definition 2.1]).
@@
Let $R$ be a Cohen-Macaulay local ring, and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then $R$ is Gorenstein in each one of the following cases.
@@
(1) The positivity of the rank of $M$ implies that $M^{\ast}\neq 0$. Since $M$ is Cohen-Macaulay and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, according to Corollary 4.4, we conclude that $M$ is totally reflexive. Consequently, $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d$. Thus, by assumption and [48, Theorem 3.1], we obtain the result.
(2) By definition, $M$ is Cohen-Macaulay, and by Theorem 4.4, $M$ is totally reflexive. Hence $R$ is Gorenstein by [11, Proposition 2.19]. ∎
Due to Proposition 6.1, is natural to ask the following.
@@
This question is a refined version of Question 4.5 of the second version of this paper, which was answered negatively in [19, Example 3.9 and Remark 3.10] by Ghosh and Samanta. The same example shows that Question 6.2 is false in general.
On the other hand, it is known that if $R$ is a local ring such that $\operatorname{G-dim}_{R}(M)<\infty$ for all $R$-modules, then $R$ is Gorenstein ([9, Theorem 1.4.9]). Thus, by considering the dual of $R$-modules it is natural to ask the following.
@@
Then motivated by Question 6.4, we obtain the following result:
@@
Let $R$ be a local ring. If $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ for all (finitely generated) $R$-modules $M$ such that $M^{\ast}\not=0,$ then $R$ is Gorenstein.
@@
Now, suppose that $\operatorname{depth}_{R}(R)\geq 1$. We may write $\mathfrak{m}=(x_{1},\ldots,x_{n})$. Let $\varphi:R\to R^{n}$ the homomorphism defined by $\varphi(r)=(rx_{1},\ldots,rx_{n})$. Since $\operatorname{depth}_{R}(R)>0$, we see that $\varphi$ is injective. Thus we have an exact sequence
$0\rightarrow R\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}R^{n}\rightarrow M\rightarrow 0$
@@
(6.1) $0\rightarrow M^{\ast}\to R^{n}\stackrel{{\scriptstyle\varphi^{\ast}}}{{\rightarrow}}R\to\operatorname{coker}(\varphi^{\ast})=k\to 0$
If $M^{\ast}=0,$ we see from (6.1) that $\operatorname{pd}_{R}(k)<\infty,$ which shows that $R$ is regular, and consequently Gorenstein. If $M^{\ast}\not=0,$ by assumption, $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, so from (6.1), $\operatorname{G-dim}_{R}(k)<\infty,$ concluding that $R$ is Gorenstein. ∎
As an immediate consequence we obtain the following corollary that answer positively Question 6.4
6.6 Corollary.
Let $R$ be a local ring such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ for all (finitely generated) $R$-modules $M$. Then $R$ is Gorenstein.
@@
Corollary 6.6 can be obtained in other ways. For instance, it could be derived from [44, Corollary 3.5]. Additionally, this corollary can be obtained as an application of Theorem 3.2. Indeed, setting $d=\dim_{R}(R)$, by the assumption and Theorem 3.2, all (finitely generated) $R$-modules satisfy the following equivalences:
@@
Then by [37, Theorem 1.4], $R$ satisfies $(G_{d})$, concluding that $R$ is Gorenstein.
@@
First, suppose $\operatorname{depth}_{R}(R)=0$. Since $\operatorname{G-dim}_{R}(\mathfrak{m}^{\ast})<\infty,$ then $\operatorname{G-dim}_{R}(\operatorname{Tr}_{R}(\mathfrak{m}))<\infty$. Since $\operatorname{depth}_{R}(R)=0$, the Auslander-Bridger formula shows that $\operatorname{Tr}_{R}(\mathfrak{m})$ is totally reflexive, and hence $\mathfrak{m}$ as well. Now, by considering the exact sequence $0\to\mathfrak{m}\to R\to k\to 0,$ we see that $\operatorname{G-dim}_{R}(k)<\infty$. Thus, we conclude that $R$ is Gorenstein.
@@
The next question that we will present is motivated by Question 4.11. Consider the following condition:
@@
Every (finitely generated) $R$-module $M$ whose dual $M^{\ast}$ is finite Gorenstein dimension, is also of finite Gorenstein dimension.
6.9 Question.
Let $R$ be a local ring. If $R$ satisfies (GDUAL), then is $R$ Gorenstein?

### v5

@@ 4. Characterization of totally reflexives
Let $R$ be a local ring of depth $t$, and let $M$ be a non-zero $R$ module. Suppose that $M$ is locally totally reflexive on the punctured spectrum of $R$. Suppose that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then the following are equivalent:
@@
This follows from Theorems 4.9 and 4.1. ∎
@@ 5. Freeness Criteria
    To discuss the Auslander-Reiten conjecture for $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$.
@@
Let $R$ be a Cohen-Macaulay local ring, and let $M$ be an $R$-module such that $\operatorname{pd}_{R}(M^{\ast})<\infty$. If $M$ is maximal Cohen-Macaulay, then $M$ is free.
Similarly to the proof of Corollary 4.4, we obtain the following.
@@
Let $R$ be a Cohen-Macaulay local ring, and let $M$ be an $R$-module such that $M^{\ast}\not=0$ and $\operatorname{pd}_{R}(M^{\ast})<\infty$. If $M$ is Cohen-Macaulay, then $M$ is free.
@@
Let $(R,\mathfrak{m},k)$ be a Cohen-Macaulay local ring. The maximal ideal $\mathfrak{m}$ is Cohen-Macaulay if and only $\operatorname{dim}(R)\leq 1$.
@@
Since $\operatorname{depth}_{R}(R)\geq 2,$ then $k^{\ast}=0=\operatorname{Ext}_{R}^{1}(k,R)$ and $\mathfrak{m}^{\ast}\cong R$. Therefore, by Corollary 5.3, $\mathfrak{m}$ is free and hence $\mathfrak{m}\cong R$. From the exact sequence (5.1), we obtain an exact sequence
@@
As $\operatorname{depth}_{R}(R)\geq 2$ and $\mathfrak{m}\cong R$, then $\operatorname{Hom}_{R}(k,\mathfrak{m})\cong\operatorname{Hom}_{R}(k,R)=0=\operatorname{Ext}_{R}^{1}(k,R)\cong\operatorname{Ext}_{R}^{1}(k,\mathfrak{m})$. Consequently, $\operatorname{Hom}_{R}(k,k)=0,$ which does not occurs.
Conversely, assume that $\operatorname{dim}(R)\leq 1$. If $\dim(R)=0$ it is clear that all $R$-modules are Cohen-Macaulay. Suppose $\operatorname{dim}(R)=1$. Since $R$ is Cohen-Macaulay, then there exists an $R$-regular element $x\in\mathfrak{m}$. Note that $x$ is also an $\mathfrak{m}$-regular element. Therefore $1\leq\operatorname{depth}_{R}(\mathfrak{m})\leq\operatorname{dim}(R)=1,$ which shows that $\mathfrak{m}$ is maximal Cohen-Macaulay. ∎
@@
Let $R$ be a local ring of depth $t\geq 1$. Let $M$ be a locally free $R$-module on the punctured spectrum such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{t-1}(M,M)=0$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t$, then $M$ is free.
@@
If $t=1$, then $\operatorname{Hom}_{R}(M,M)\cong\operatorname{Ext}_{R}^{0}(M,M)=0$, whence $M$ is zero and hence free. Suppose that $t\geq 2$. By Theorem 4.1, $M$ is totally reflexive and hence $\operatorname{Tr}(M)$ as well. Thus, $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{i}(\operatorname{Tr}(M),R)=0$ for all $i>0$. Therefore, the result follows from [AuslanderReitenConjectureforNormalRings, Proposition 2.8]. ∎
@@
Let $R$ be a local ring of depth $t\geq 1$. Let $M$ be a locally free $R$-module on the punctured spectrum such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $0pt(M)\geq t$ and $\operatorname{Ext}_{R}^{t-1}(M,M)=0$, then $M$ is free.
@@
Let $R$ be a ring satisfying $(S_{1})$ and $X$ be a subset of $\operatorname{Spec}(R)$ containing $X^{1}(R)$. Let
[…]
Let $M$ be a locally free $R$-module on $X$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. If $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $s-1\leq j\leq t-1$, then $M$ is projective.
@@
We may suppose that $X\not=\operatorname{Spec}(R)$. We show that $M_{\mathfrak{p}}$ is free for all $\mathfrak{p}\in\operatorname{Spec}(R)$. By assumption this is true if $\mathfrak{p}\in X$. So, suppose that $\X\mathfrak{p}\in\operatorname{Spec}(R)\backslash$ X. Therefore $\operatorname{ht}(\mathfrak{p})\geq 2$ and $\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\geq 1$ since $R$ satisfies $(S_{1})$. By definition of $s$ and $t$, note that $s\leq\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\leq t$, so $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(M_{\mathfrak{p}},R_{\mathfrak{p}})=0$ for all $1\leq i\leq\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})$ and $\operatorname{Ext}_{R_{\mathfrak{p}}}^{\operatorname{depth}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})-1}(M_{\mathfrak{p}},M_{\mathfrak{p}})=0$. Hence, by Proposition 5.5, $M_{\mathfrak{p}}$ is free. ∎
@@
([AuslanderReitenConjectureforNormalRings, Remark 2.11]) Let $R$ be a Cohen-Macaulay local ring of dimension $d$ and $n$ be an integer such that $1\leq n\leq d-1$. Then
@@
[…]
@@
(1) We claim that $M$ is locally free on $X^{n}(R)$. In fact, given $\mathfrak{p}\in X^{n}(R)$, we have by hypothesis that $\operatorname{pd}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})<\infty$ and $\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}(M_{\mathfrak{p}},R_{\mathfrak{p}})=0$ for all $1\leq i\leq d$. Hence, by [CommutativeRingTheory, p. 154, Lemma 1(iii)], $M_{\mathfrak{p}}$ is free as $R_{\mathfrak{p}}$-module. Thus item (1) follows immediately from Theorem 5.7 and Remark 5.8.
(2) Let $p=\operatorname{G-dim}_{R}(M)$. We prove the result by considering two cases.
@@
(2.b) Suppose $p>0$. We have $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i\geq p+1$. By [AuslanderReitenConjectureforNormalRings, Lemma 2.6], we get
@@
Let $R$ be a Cohen-Macaulay normal local ring of dimension $d\geq 2$, and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Suppose that $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq d$ and $1\leq j\leq d-1$. Then $M$ is free.
@@
In [FiniteHomologicalDimensionOfHomv3, Corollary 6.9(2)], Dey and Ghosh demonstrated that the Auslander-Reiten conjecture holds for (finitely generated) $R$-modules $M$ satisfying $\operatorname{G-dim}_{R}(M)<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$. Motivated by this, we investigate whether the conjecture remains valid when we consider $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ instead of $\operatorname{G-dim}_{R}(M)<\infty$ in the referenced case.
@@
    $\operatorname{Hom}_{R}(M,M)$ is free and $\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq j\leq t-1$.
@@
    $\operatorname{Hom}_{R}(M,M)$ has finite projective dimension, and $\operatorname{Ext}_{R}^{i}(M,R)=\operatorname{Ext}_{R}^{j}(M,M)=0$ for all $1\leq i\leq t$ and $1\leq j\leq t-1$.
@@
$(2)\Rightarrow(3)$. We only need to prove that $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq t$. For this, we will show that $M$ is totally reflexive. In view of [FiniteHomologicalDimensionOfHomv3, Corollary 7.6(2)], $M$ satisfies $(\widetilde{S}_{t})$. Therefore, $M$ is totally reflexive by Theorem 4.1.
$(3)\Rightarrow(1)$. By Theorem 4.1, $M$ is totally reflexive. Therefore by [FiniteHomologicalDimensionOfHomv3, Corollary 6.8], $M$ is free. ∎
@@
Let $R$ be a ring. The Auslander-Reiten conjecture holds true for all (finitely generated) $R$-modules $M$ such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(M,M))<\infty$.
@@
Let $R$ be a local ring and let $C$ be a semidualizing $R$-module. If $\operatorname{G-dim}_{R}(C^{\ast})<\infty$, then $C\cong R$.
@@ 6. Gorenstein criteria and related questions
Throughout this section, let $(R,\mathfrak{m},k)$ be a local ring. In this section, we aim to provide a number of criteria for a local ring $R$ to be Gorenstein in terms of the dual of certain modules having finite Gorenstein dimension. In this sense, we discuss some related questions to this subject.
The first criterion that we present in this section is an application of Corollary 4.4. Before stating it, first recall that if $R$ is a ring with total quotient ring $Q$, then an $R$-module $M$ is said to have a (generic) rank, denoted by $\operatorname{rank}(M)$, if $M\otimes_{R}Q$ is a free $Q$-module of rank $\operatorname{rank}(M)$. Let $e(R)$ denote the Hilbert-Samuel multiplicity of $R$ and $\mu(M)$ denote the minimum number of generators of $M$. Moreover, if $M$ is Cohen-Macaulay and $e(M)=\mu(M)$, then $M$ it is said to be an Ulrich module ([AlmostGorensteinRingsTowardsATheoryOfHigherDimension, Definition 2.1]).
@@
Let $R$ be a Cohen-Macaulay local ring, and let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(M^{\ast})<\infty$. Then $R$ is Gorenstein in each one of the following cases:
@@
(1) The positivity of the rank of $M$ implies that $M^{\ast}\neq 0$.
Since $M$ is Cohen-Macaulay and $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, according to Corollary 4.4, we conclude that $M$ is totally reflexive. Consequently, $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq d$. Thus, by assumption and [Ulrich, Theorem 3.1], we obtain the result.
(2) By definition, $M$ is Cohen-Macaulay, and by Corollary 4.4, $M$ is totally reflexive. Hence $R$ is Gorenstein by [ComplexityandrigidityofUlrichModules, Proposition 2.19]. ∎
Due to Proposition 6.1, it is natural to ask the following.
@@
This question is a refined version of Question 4.5 of the second version of this paper, which was answered negatively in [ARCconjectureForModulesWhoseSelfDualHasFiniteCompleteIntersectionDimension, Example 3.10 and Remark 3.11] by Ghosh and Samanta. The same example shows that Question 6.2 is false in general.
On the other hand, it is known that if $R$ is a local ring such that $\operatorname{G-dim}_{R}(M)<\infty$ for all $R$-modules, then $R$ is Gorenstein ([GorensteinDimensions, Theorem 1.4.9]). Thus, by considering the dual of $R$-modules it is natural to ask the following.
@@
Then motivated by Question 6.4, we obtain the following result which says that its answer is positive.
@@
Let $R$ be a local ring. If $\operatorname{G-dim}_{R}(M^{\ast})<\infty$ for all (finitely generated) $R$-modules $M$, then $R$ is Gorenstein.
@@
Now, suppose that $\operatorname{depth}_{R}(R)\geq 1$. We may write $\mathfrak{m}=(x_{1},\ldots,x_{n})$. Let $\varphi:R\to R^{n}$ be the homomorphism defined by $\varphi(r)=(rx_{1},\ldots,rx_{n})$. Since $\operatorname{depth}_{R}(R)>0$, we see that $\varphi$ is injective. Thus we have an exact sequence
$0\rightarrow R\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}R^{n}\rightarrow M\rightarrow 0.$
@@
(6.1) $0\rightarrow M^{\ast}\to R^{n}\stackrel{{\scriptstyle\varphi^{\ast}}}{{\rightarrow}}R\to\operatorname{coker}(\varphi^{\ast})=k\to 0.$
Since $\operatorname{G-dim}_{R}(M^{\ast})<\infty$, from (6.1), we see that $\operatorname{G-dim}_{R}(k)<\infty,$ concluding that $R$ is Gorenstein. ∎
@@
Theorem 6.5 can be obtained in other ways. For instance, it could be derived from [ktorsionlessmoduleswithfiniteGorensteindimension, Corollary 3.5]. Additionally, the theorem can be obtained as an application of Theorem 3.2. Indeed, setting $d=\dim_{R}(R)$, by the assumption and Theorem 3.2, all (finitely generated) $R$-modules satisfy the following equivalences:
@@
Then by [WhenArenSysymodulesnTorsionfree, Theorem 1.4], $R$ satisfies $(G_{d})$, concluding that $R$ is Gorenstein.
@@
First, suppose $\operatorname{depth}_{R}(R)=0$. Since $\operatorname{G-dim}_{R}(\mathfrak{m}^{\ast})<\infty,$ then $\operatorname{G-dim}_{R}(\operatorname{Tr}(\mathfrak{m}))<\infty$. Since $\operatorname{depth}_{R}(R)=0$, the Auslander-Bridger formula shows that $\operatorname{Tr}(\mathfrak{m})$ is totally reflexive and hence $\mathfrak{m}$ as well. Now, by considering the exact sequence $0\to\mathfrak{m}\to R\to k\to 0,$ we see that $\operatorname{G-dim}_{R}(k)<\infty$. Thus, we conclude that $R$ is Gorenstein.
@@
Motivated by Question 4.11, we consider the following condition:
@@
Every (finitely generated) $R$-module $M$ whose dual $M^{\ast}$ is of finite Gorenstein dimension, is also of finite Gorenstein dimension.